# references.bib

### v3

@@ Abstract
Remote sensing map products are used to obtain estimates of environmental quantities, such as deforested area or the effect of conservation zones on deforestation. However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize. Thus, population-level estimates from such maps may be biased. In this paper, we compare several uncertainty quantification methods — stratified estimator, post-stratified estimator, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate unbiased estimates. Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product. Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling. This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps. Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 1 Introduction
Remote sensing maps are widely used to obtain estimates for environmental quantities [hansen2013high, baccini2012estimated, potapov2022global, wang2019great, deines2019satellites]. Map products created by training machine learning models on remotely sensed data allow researchers to cheaply access predictions for a variable of interest at a high resolution over large regions.
@@
Additionally, errors in map products are often difficult to characterize due to the complexity of input data and lack of interpretability in the ML models used to generate the maps. Examples of systematic errors in map products trained on satellite data include “nonrandom misclassification, saturation effects, atmospheric effects, and cloud cover,” all of which may lead to biased inference [jain2020benefits]. This has prevented environmental scientists, economists, and policymakers from using remote sensing maps in their analyses. The World Bank has cited uncertainty in remote sensing maps as a barrier to their use in water resources management [world2018mainstreaming], and the United Nations Task Team on Satellite Imagery and Geospatial Data wrote that the “shift in paradigm from traditional statistical methods (e.g., counting, measuring by humans) towards estimation from sensors (with associated aspects of validity, approximation, uncertainty)… will require convincing, statistically sound results” [united2017earth]. Correctly accounting for map errors, then, is crucial for maximizing the potential of remote sensing data to expand scientific knowledge and advance practical operations.
@@
For those who currently reject estimation from remote sensing as too unreliable, the alternative is to use ground truth observations for the variable of interest. However, ground truth data is often sparse because it is expensive to collect, as it may require fieldwork or manual labelling of remote sensing imagery. Few ground truth points result in greater uncertainty (e.g., wider confidence intervals) in estimates. Another alternative is to use both ground truth data and remote sensing maps to generate estimates, using the ground truth points to correct for biases that may exist in the map products. This generally reduces uncertainty in the estimates compared to using ground truth alone, since map products contain a large amount of data that, while possibly biased, still give useful information about the true values we wish to measure.
Figure 1: Motivating example: map-only estimate is biased. The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points. The map-only slope coefficient estimate falls outside the confidence intervals for all methods that use ground truth points.
@@
In this work, we discuss methods that correct map products with ground-truth data to give unbiased confidence intervals. We give a brief preview here. We seek to estimate logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates, including terrain slope. Using nearly one million remote sensing map product points and $1386$ ground truth points, we compare the $95\%$ confidence intervals for the slope coefficient with four methods (Figure 1) described shortly. See Section 4.1 for a complete description of the experimental setup.
The “classical” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products. Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being unbiased.
Notably, the slope coefficient computed from the map product alone is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points. That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation. Methods like PPI allow us to use ground truth points to correct for bias in the map product and thus avoid misleading conclusions.
Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products. Stratified, post-stratified, and prediction powered inference estimators use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest. We also compare these methods against classical uncertainty quantification (using only ground truth) and map-only estimators.
@@
In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate unbiased estimates with low uncertainty (Figure 2). We also compare these methods against the classical estimator (using only ground truth) and map-only estimator (using only a map product). We consider the following three methods:
  • Stratified estimator uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates. This method has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good].
  • Post-stratified area estimator uses the confusion matrix between ground truth and map product land cover labels to correct for misclassification bias in the map product, when estimating the area of a land cover class. This method is also commonly used in the remote sensing community [card1982using, stehman2013estimating, olofsson2013making].
  • Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone.
We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation. Area estimation includes estimating both total area occupied by a land cover/land use category and total change in land cover/land use over time. The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock. A map-only estimator for area is known as a pixel-counting estimator. Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables. For regressions, this paper only accounts for uncertainty in the remotely sensed dependent variable, while the covariates are assumed to have no errors. To illustrate the methods, we consider four specific remote sensing estimation tasks spanning crop area estimation, deforestation quantification, and associating deforestation and tree cover to other geophysical variables; see Table 1. Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
Estimation task Target variable Region Year Map product(s) Ground truth Covariates Link to code
Area Maize Iowa 2022 USDA NASS Cropland Data Layers (CDL) [nass2024usda] Simulated N/A link
Area Deforestation Brazilian Amazon 2000-2015 1. NASA Global Forest Cover Change [townshend2016global]
2. Amazon ecoregion
deforestation strata [bullock2020satellite]
$n=1386$
@@ 2. Amazon ecoregion
N/A link
Logistic
regression
coefficients
Distance from roads, distance from rivers, elevation, slope link
Linear
regression
coefficients
Tree cover Contiguous USA 2021 USFS Tree Canopy Cover [forest2023usfs] $n=983$
manually labelled Google Maps high-resolution satellite images
Aridity index,
@@
Table 1: Overview of use cases and datasets. We estimate maize area in Iowa, deforestation area and logistic regression coefficients in the Brazilian Amazon, and tree cover linear regression coefficients in the United States. The Iowa maize area use case is a simulation: we treat a small random sample of CDL map product points as “ground truth” data and observe the effects of artificially degrading the map product relative to the ground truth. Code implementations of the use cases are linked in the rightmost column.
Combining ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than classical estimates. As we will show, map-only estimators without bias correction can result in completely incorrect area and regression coefficient estimates. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a (simple or stratified) randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
@@
The remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimates. This estimator uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product [card1982using, stehman2013estimating, olofsson2013making]. This method only applies to area estimation, not regression coefficient estimation tasks. One contribution of the present work is a comparison of the post-stratified area estimator and PPI. We find both experimentally and mathematically that these two methods produce similar outputs for area estimation. For other methods, [gallego2004remote] provides a survey of area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators.
@@
Methods have also been developed to reduce bias in regression coefficient estimates from remote sensing data. For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico. They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification error as dependent on environmental covariates via a logit model. Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa. They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover. Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US. They use a linear measurement error model and apply multiple imputation: a statistical method that uses a small amount of ground truth data to estimate remote sensing measurement error and impute corrected values for remotely-sensed variables. Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
@@
Another approach to uncertainty quantification is conformal prediction, which generates confidence sets for machine learning predictions at individual data points. For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time). This approach differs from our work because our methods estimate total area and regression coefficients over an entire region, rather than quantifying uncertainty at individual points or pixels.
@@
Unlike previous works, this paper is the first work to apply prediction-powered inference to remote sensing use cases in area and regression coefficient estimation. Unlike stratified random sampling, PPI does not require that ground truth points be sampled according to a pre-determined stratification map; as a result, it allows for greater flexibility in ground truth data usage across different downstream applications. Additionally, this is the first work to combine map product and ground truth data to perform uncertainty quantification on remote sensing regression coefficients without assuming an error structure on the map product predictions.
2 Background: Uncertainty Quantification
We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables. As two baselines, we first consider methods that use only the ground-truth data (which is unbiased but loses information) or only use the remote-sensing data (which may be biased). We then turn to methods that use both sources of data to produce unbiased confidence intervals. See Table 2 for a summary.
Method Data Used Unbiased? Caveats
Map Ground Truth
Map-only
Classical
Stratified Ground truth points sampled
using stratified map
Post-stratified Area estimation only
PPI
Table 2: Summary of uncertainty quantification methods. PPI is the most flexible method that incorporates both ground truth and map product data.
2.1 Baseline methods
Classical: use only the ground-truth data.
First, classical uncertainty quantification uses ground truth observations, sampled uniformly at random, to estimate the quantity of interest. For example, suppose there is a target variable $Y$ that can be observed over a given area, and we wish to generate a 95% confidence interval for its population mean $\theta^{\star}=E[Y]$. If we have $n$ ground truth observations $Y_{1},Y_{2},\dots Y_{n}$ sampled uniformly at random, we can compute the sample mean
$\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$
and standard error
$SE=\sqrt{\frac{1}{n}\mathrm{Var}(Y_{i})}=\frac{1}{n}\sqrt{\sum_{i=1}^{n}(Y_{i}-\hat{\theta})^{2}}$
to produce the classical 95% confidence interval
$C=(\hat{\theta}-1.96\cdot SE,\ \hat{\theta}+1.96\cdot SE).$
This method is conservative — if the size $n$ of the ground truth dataset is small, the classical confidence interval will be large.

### v4

# Regression coefficient estimation from remote sensing maps

@@ Abstract
Remote sensing map products are used to estimate regression coefficients relating environmental variables, such as the effect of conservation zones on deforestation. However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize. Thus, population-level estimates from such maps may be biased. In this paper, we apply prediction-powered inference (PPI) to regression coefficient estimation. PPI generates unbiased estimates by using a small amount of randomly sampled ground truth data to correct for bias in large-scale remote sensing map products. Applying PPI across multiple remote sensing use cases in regression coefficient estimation, we find that it results in estimates that are (1) more reliable than using the map product as if it were 100% accurate and (2) have lower uncertainty than using only the ground truth and ignoring the map product. Empirically, we observe effective sample size increases of up to 17-fold using PPI compared to only using ground truth data. This is the first work to estimate remote sensing regression coefficients without assumptions on the structure of map product errors. Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 1 Introduction
Remote sensing maps are widely used to obtain estimates for environmental quantities, such as cropland expansion over time or the effect of protected areas on deforestation [hansen2013high, baccini2012estimated, potapov2022global, wang2019great, deines2019satellites]. Map products created by training machine learning models on remotely sensed data allow researchers to cheaply access predictions for a variable of interest at a high resolution over large regions.
@@
Errors in map products are often difficult to characterize due to the complexity of input data and lack of interpretability in the ML models used to generate the maps. Examples of systematic errors in map products trained on satellite data include “nonrandom misclassification, saturation effects, atmospheric effects, and cloud cover,” all of which may lead to biased inference [jain2020benefits]. This has prevented environmental scientists, economists, and policymakers from using remote sensing maps in their analyses. The World Bank has cited uncertainty in remote sensing maps as a barrier to their use in water resources management [world2018mainstreaming], and the United Nations Task Team on Satellite Imagery and Geospatial Data wrote that the “shift in paradigm from traditional statistical methods (e.g., counting, measuring by humans) towards estimation from sensors (with associated aspects of validity, approximation, uncertainty)… will require convincing, statistically sound results” [united2017earth]. Correctly accounting for map errors, then, is crucial for maximizing the potential of remote sensing data to expand scientific knowledge and advance practical operations.
@@
For those who currently reject estimation from remote sensing as too unreliable, the alternative is to use ground truth observations for the variable of interest. However, ground truth data is often sparse because it is expensive to collect, as it may require fieldwork or manual labelling of remote sensing imagery. Few ground truth points result in greater uncertainty (i.e., wider confidence intervals) in estimates.
Another alternative is to use both ground truth data and remote sensing maps to generate estimates, using the ground truth points to correct for biases that may exist in the map products. This generally reduces uncertainty in the estimates compared to using ground truth alone, since map products contain a large amount of data that, while possibly biased, still give useful information about the true values we wish to measure.
As an example of this approach, the remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimation. A pixel-counting estimator, which uses only a map product to estimate area of a land cover class, may be biased due to misclassification error in the map product [olofsson2020mitigating]. An unbiased area estimate can be obtained by using randomly sampled ground truth data [olofsson2014good, stehman2019key], but if ground truth observations are sparse, this may result in less precise estimates (wider confidence intervals) than desired. To resolve this, the post-stratified area estimator [card1982using, stehman2013estimating, olofsson2013making] combines a map product with a small amount of ground truth data, using the confusion matrix between land cover ground truth and map product points to correct for bias in the map product. By leveraging the map product, the post-stratified estimator also produces narrower confidence intervals compared to only using ground truth data. Other area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators, are described in [gallego2004remote].
Figure 1: Overview of estimation methods using ground truth data and remote sensing map products. The post-stratified estimator (for area estimation) and prediction-powered inference (for regression coefficient estimation) use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest. In this paper, we focus on regression coefficient estimation using PPI. We also compare PPI against GT-only (using only ground truth) and map-only estimators.
The remote sensing community also uses map products for regression coefficient estimation: estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables. Unlike area estimation with the post-stratified estimator, regression coefficient estimation from remote sensing maps still suffers from bias due to map product errors. An unbiased estimate can be obtained by running a regression on randomly sampled ground truth data [sarndal2003model], but similarly to the area estimation case, this may result in less precise estimates than desired.
In this paper, we address this problem by applying prediction-powered inference (PPI) to regression coefficient estimation (Figure 1). PPI [angelopoulos2023prediction, angelopoulos2023ppi++, Kluger25GeneralizingPPI] is a recently developed unbiased estimation method that uses a small set of randomly sampled ground truth observations ($n$ points) to calibrate and correct for bias in a much larger set of machine learning-generated predictions ($N$ points, where $N$ is much larger than $n$). This method also results in more precise estimates (higher statistical power, smaller confidence intervals) than using the ground truth points alone. We apply PPI to regression use cases where the machine learning predictions are obtained from large-scale remote sensing map products. We compare results with the ground truth-only (GT-only) estimator (using only ground truth points) and map-only estimator (using only a map product). These estimators are described in detail in Section 2.
Figure 2: Motivating example: PPI produces unbiased coefficient estimate with lower variance than using ground truth alone. We estimate the linear regression coefficient relating US forest cover to elevation (with $N=67968$ remote sensing map points and $n=500$ ground truth points). Left: Map-only, ground truth-only, and PPI point estimates and 95% confidence intervals. Right: Histogram of PPI and GT-only point estimates from repeating the experiment with 100 random sets of $n=500$ ground truth points.
@@
We give a brief preview of how we use PPI to correct map products with ground-truth data to obtain an unbiased regression coefficient estimate.
We seek to estimate linear regression coefficients for the association between US forest cover and a few covariates, including elevation. We use $N=67968$ remote sensing map product points and $n=500$ ground truth points. See Section 3.1 for a complete description of the experimental setup.
We compare the $95\%$ confidence intervals for the elevation coefficient estimates from PPI, GT-only, and map-only methods (Figure 2, left panel). The “GT-only” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products. Meanwhile, PPI uses the map product along with the ground truth points to produce a much smaller confidence interval while still being unbiased. The PPI and GT-only confidence intervals both contain the true coefficient, while the confidence interval estimated from the map product alone does not. That is, if we had used only the map product, we would have underestimated the magnitude of the negative relationship between forest cover and elevation. PPI avoids misleading conclusions by using ground truth points to correct for bias, and at the same time, it gives smaller confidence intervals than using only the ground truth points.
We repeat the experiment using the same set of $N=67968$ map product points and 100 different randomly chosen sets of $n=500$ ground truth points (see Appendix C.4 for more detail). A histogram of the 100 PPI and GT-only elevation coefficient point estimates is shown in the right panel of Figure 2. As before, we see that the PPI and GT-only estimators are both unbiased relative to the true coefficient, but the PPI estimates are more precise as they have a lower variance than the GT-only estimates.
@@
In this paper, we illustrate the use of PPI on four specific regression coefficient estimation tasks associating deforestation, tree cover, housing price, and forest cover to other geospatial variables; see Table 1. We then simulate map error to understand its influence on the bias and variance of map-only, GT-only, and PPI estimators. We demonstrate that PPI can be applied when there is error in the remotely sensed response variable, the remotely sensed covariates, or both. Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
Regression type Response variable Covariates Region Year Map product(s) Ground truth Link to code
Logistic Deforestation Distance from roads, distance from rivers, elevation, slope Brazilian Amazon 2000-2015 NASA Global Forest Cover Change [townshend2016global] $n=1386$
@@
link
Linear Tree cover Aridity index,
@@
Contiguous USA 2021 USFS Tree Canopy Cover [forest2023usfs] $n=983$
manually labeled images (Sec. 3.1)
@@
Linear Housing price Income, nightlights, road length Contiguous USA 2010-2021 MOSAIKS [rolf2021generalizable] $n=500$
from [rolf2021generalizable]
link
Linear Forest cover Elevation,
population
Contiguous USA 2010 MOSAIKS [rolf2021generalizable] $n=500$
from [rolf2021generalizable]
link
Table 1: Overview of use cases and datasets. To illustrate the mechanics and advantages of PPI, we estimate linear regression coefficients for tree cover, housing price, and forest cover in the United States. Code implementations of the use cases are linked in the rightmost column.
Using PPI to combine ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than GT-only estimates. As we will show, map-only estimators without bias correction can result in incorrect regression coefficient estimates.
To improve the utility of machine learning-generated remote sensing maps for downstream regression applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in PPI alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data). This holdout set can be obtained by simple (uniform) random sampling or other sampling designs such as stratified, weighted, or clustered random sampling [Kluger25GeneralizingPPI].
@@
Previous works have developed methods to reduce bias in regression coefficient estimates from remote sensing data. For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico. They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification probabilities as a parametric function of environmental covariates. However, this approach imposes assumptions about the structure of the misclassification errors and moreover is only developed for settings with binary response variables. Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa. They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover; however, they do not provide an approach for constructing confidence intervals. Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US. To remedy this issue, they propose using a Bayesian linear measurement error model and applying multiple imputation: a statistical method that involves estimating the missing values in a dataset and their uncertainties and fitting the regression on multiple randomly imputed versions of the dataset. Note that this method requires an assumption about the error structure of the remotely-sensed variables and does not have any frequentist coverage guarantees. For example, in a subset of the simulation-based experiments in [proctor2023parameter] 95% confidence intervals from multiple imputation only contained the true value of the regression coefficient in fewer than 80% of the simulations as opposed to the desired $\sim$95of the simulations. In contrast to existing approaches in the remote sensing literature, PPI does not make any modelling assumptions about the measurement error structure, and under certain regularity conditions, its confidence intervals are guaranteed to attain the desired coverage [angelopoulos2023prediction, angelopoulos2023ppi++, Kluger25GeneralizingPPI].
@@
Another approach to uncertainty quantification is conformal prediction [vovk2005algorithmic, angelopoulos2023conformal], which generates confidence sets for machine learning predictions at individual data points. For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time). This approach differs from our work because PPI estimates regression coefficients over an entire region, rather than quantifying uncertainty of map estimates at individual points or pixels.
@@
Unlike previous works, this paper is the first work to apply prediction-powered inference to remote sensing use cases in regression coefficient estimation. This is also the first work to combine map product and ground truth data to estimate remote sensing regression coefficients without assuming an error structure on the map product predictions.
2 Methods: Prediction-Powered Inference for Regression
Regressions are commonly used in the environmental sciences and economics to identify relationships — associative or causal — between covariates and a binary response (logistic regression) or between covariates and a continuous response (linear regression). Logistic regression coefficients give the change in the log-odds ratio of a binary response corresponding to a one-unit increase in a covariate, when all other covariates are held constant. Linear regression coefficients give the change in the value of a continuous response corresponding to a one-unit increase in a covariate, when all other covariates are held constant. Thus, we can interpret these coefficients as effect sizes of the covariates.
Historically, scientists would run regressions using ground truth environmental samples, but regressions on data from remote sensing maps are becoming increasingly common, e.g., [taylor2023fertilizer, deines2023recent, he2022marked, taylor2022wetlands, deines2019satellites, lobell2022globally]. When covariate or response variable values are machine learning predictions from a map product, errors in the machine learning model may result in biased regression coefficients (as we have seen in Figure 2).
In this section, we formally describe how to use prediction-powered inference (PPI) to generate regression coefficient estimates and confidence intervals when the response variable $Y$ and/or covariates $X$ are remotely sensed map products. As two baselines, we consider the ground truth-only estimator that only uses ground-truth data (which is unbiased but does not leverage remote sensing maps) and the map-only estimator that only uses remote sensing maps (which may be biased). We then describe PPI, which uses both sources of data to produce unbiased confidence intervals. Figure 1 summarizes the methods. As a technical note, this paper focuses on inference in settings where both the ground truth data and the larger sample from a remote sensing map product are assumed to be randomly sampled from a superpopulation, as opposed to a design-based inference framework for reasons discussed in Appendix A.
Suppose we have a remote sensing data product with a large number $N$ of predictions $(\hat{X}^{\prime}_{1},\hat{Y}^{\prime}_{1})$, $(\hat{X}^{\prime}_{2},\hat{Y}^{\prime}_{2})$, $\dots$, $(\hat{X}^{\prime}_{N},\hat{Y}^{\prime}_{N})$ where $X\in\mathbb{R}^{p}$ is a vector of covariates and $Y\in\mathbb{R}$ is the response variable. Suppose that for a small calibration set of $n\ll N$ points sampled uniformly at random, we have ground truth labels $(X_{1},Y_{1})$, $(X_{2},Y_{2})$$\dots$, $(X_{n},Y_{n})$ with corresponding remote sensing predictions $(\hat{X}_{1},\hat{Y}_{1}),(\hat{X}_{2},\hat{Y}_{2}),\dots,(\hat{X}_{n},\hat{Y}_{n})$. (Note that in some use cases, ground truth data may be widely available either for the response variable, for a subset of the covariates, or for a subset of both. In these cases, we would use ground truth instead of remotely sensed values for the corresponding components of the $(\hat{X},\hat{Y})$ and $(\hat{X}^{\prime},\hat{Y}^{\prime})$ vectors.)
2.1 Linear regression
In the case of linear regression, we wish to estimate 95% confidence intervals for the regression coefficients $\beta_{*}$ giving the best fit to the linear model
$Y=\beta^{\mathsf{T}}X+\varepsilon,$
where $\varepsilon$ is mean 00 and uncorrelated with $X$. (We assume that the first element of the vector $X$ is a constant 1, so that the first regression coefficient $\beta_{0}$ is an intercept term.)
Ground truth-only estimator.
The ground truth-only (GT-only) estimator uses only the $n$ ground truth calibration datapoints to compute the regression coefficient estimate
$\hat{\beta}_{\text{calib}}=\operatorname*{arg\,min}_{\beta}\sum_{i=1}^{n}(Y_{i}-\beta^{\mathsf{T}}X_{i})^{2}.$
However, the standard error of this ground truth coefficient estimate will be large when the number of ground truth observations $n$ is small, which is often the case in environmental applications. As a result, the GT-only 95% confidence interval will be large.